OpenStudy (anonymous):

help 7-(-21) and can you tell me how you got the answer? thank you

OpenStudy (jdoe0001):

what do you mean?

OpenStudy (anonymous):

how do u find the answer? is it a negative times a negative.

OpenStudy (jdoe0001):

yes, signs rules if the signs are the equal, their product is " + " if the signs are different, their product is " - " \(\bf - \times - = +\\ 7-(-21) \implies 7-1(-21) \implies -1 \times -21 = +21\\ 7-(-21) \implies 7 + 21\)

OpenStudy (anonymous):

so the answer would be 28 correct?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thank you soooooo munch

OpenStudy (jdoe0001):

yw